OpenStudy (anonymous):

how can you solve this, without using a calculator Log (0.5) = x I know the answer is -0.3, but I would like to know how to do it without using a calculator ( since you can not convert 0.5 into a power of 10 )

Nnesha (nnesha):

log have base 10 \[\huge\rm log_{10} .5 =x\] change this to exponential form

OpenStudy (anonymous):

5 x 10^-1

OpenStudy (anonymous):

yep and then?

Nnesha (nnesha):

10^-1 ?

OpenStudy (anonymous):

lol what ?

OpenStudy (anonymous):

0.5 is 5 x 10^1 is the power of 10 of 0.5 yeah, but since its 5 x 10^-1, and not 1.0 x 10^(- something), how can u do it by hand?

Nnesha (nnesha):

how did you get 10^-1 now i'm confused >.<  isn't that suppose to loolk like this

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

oh, no i was just re writing the equation as log (5x10^-1) = x

OpenStudy (jdoe0001):

@javawarrior what does "solve" mean in this case? is not -3 btw

OpenStudy (anonymous):

so what's the next step ?

OpenStudy (jdoe0001):

or -0.3 either, since it gives a float

OpenStudy (anonymous):

just solving for x

Nnesha (nnesha):

answer is -.301

OpenStudy (anonymous):

yeah thats the answer, but I want to know how you could do it without using a calculator

OpenStudy (jdoe0001):

solving for "x"? well, is sorta already done then log(0.5) = x <----- is already solved, nothing to do further I'd say

Nnesha (nnesha):

\[\frac{ \log .5 }{ \log 10 }\] that's who we get -.30 but she wants to know who t get answer without using calc ~.^

OpenStudy (anonymous):

there's a exam question: working out the pH of a buffer, using the henderson-hassleback equation right, and the exam doesn't permit calculators.

OpenStudy (anonymous):

so if they reckon you can do it without a calculator, was wondering how you could it lol

OpenStudy (jdoe0001):

well..... the number I get is a float so.... unless you'd end up with a fraction, you'd need a calculator or take a good while getting -0.30102999566398119521

OpenStudy (anonymous):

the whole equation is 7.21 + log(0.5) = x

OpenStudy (anonymous):

oh they said in the question that log 2 = 0.30

OpenStudy (jdoe0001):

maybe you're expected to leave it in log terms

OpenStudy (jdoe0001):

heheh

Nnesha (nnesha):

hmmmm

OpenStudy (anonymous):

is there anyway you can relate log 2 = 0.30 with log 0.5 = x ?

OpenStudy (jdoe0001):

so... come on now... as paul harvey would say "what's the rest of the story"?

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

hahaha sorry guys, just saw it

OpenStudy (jtvatsim):

Here is an interesting site: http://www.science-site.net/logcalc.htm

OpenStudy (jdoe0001):

this sounds like what's the rabbit color? ohh and the the rabbit is albino ohh and the rabbit has a ferret parent ohh and the rabbit is aquatic.. and so on

OpenStudy (anonymous):

thanks jvatsim, going to start reading it now

OpenStudy (anonymous):

so none of you guys have a clue ?

OpenStudy (jdoe0001):

@javawarrior give us the full story, rather than the remnant ad yes, I do see the relevance of log(2)

OpenStudy (jtvatsim):

So if I am reading this right... the procedure is as follows: log(0.5) = log(5 x 10^-1) = log5 + log(10^-1) = log 5 - 1. Then log(5) = log(10/2) = log 10 - log 2 = 1 - 0.30 = 0.70. So, log 5 - 1 = 0.70 - 1 = -0.30

OpenStudy (anonymous):

solve for x 7.21 + log0.5 = x where log2=0.30

OpenStudy (jdoe0001):

or post a quick screenshot of the material, so we could see what you've got

OpenStudy (jdoe0001):

k

OpenStudy (anonymous):

I think jvatsim has got it, but is there a quicker way ? probably not right ?

OpenStudy (anonymous):

thanks @jtvatsim

OpenStudy (jtvatsim):

quicker way... memorize all tables of logarithms... :) but not practical

OpenStudy (jdoe0001):

\(\bf 7.21 + log(0.5) = x\implies 7.21+log\left( \cfrac{1}{2}\right)=x \\ \quad \\ {\color{brown}{ log(a)-log(b)\to log\left( \frac{a}{b} \right) }} \\ \quad \\ 7.21+log\left( \cfrac{1}{2}\right)=x\implies 7.21+[log(1)-log(2)]=x \\ \quad \\ 7.21+[log_{10}(1)-log_{10}(2)]=x \\ \quad \\ {\color{brown}{ log_{10}(1)\to \square \iff 10^\square =1\implies \square =0 }} \\ \quad \\ 7.21+[{\color{brown}{ 0}}-log_{10}(2)]=x\implies 7.21-log(2)=x\)

OpenStudy (michele_laino):

hint: \[\log \left( {0.5} \right) = \log \left( {\frac{1}{2}} \right) = - \log 2\]